what is the parabola graph of 6x^2=5y
Answer by anantha(86)

sol:
6x^2=5y
y=6x^2/5
now put x=1 then y=6*(1^2)/5=6/5=1.2 here(x,y)=(1,1.2)
now put x=2 then y=6*(2^2)/5=24/5=4.8 here (x,y)=(2,4.5)
now put x=3 then y=6*(3^2)/5=54/5=10.8 here(x,y)=(3,10.8)
now put x=0 then y=6*(0^2)/5=0/5=0 here (x,y)=(0,0)
now put x=-1 then y-=6*(-1^2)/5=6/5=1.2 here (x,y)=(-1,1.2)
now put x=-2 then y=6*(-2^2)/5=24/5=4.8 here (x,y)=(-2,4.8)
now put x=-3 then y=6*(-3^2)/5=54/5=10.8 here (x,y)=(-3,10.8)
the points are (1,1.2),(2,4.5),(3,10.8),(0,0),(-1,1.2),(-2,4.5),(-3,10.8)
using these points draw the graph